OpenStudy (anonymous):

A line in the Paschen series of hydrogen has a wavelength of 1280 nm. From what state did the electron originate?

OpenStudy (jfraser):

plug into the rydberg equation, solving for \(n_1\)\[\frac{1}{\lambda} = R(\frac{1}{n_1^2} - \frac{1}{n_2^2})\]